OpenStudy (anonymous):

\[\log _{10}x+\log _{10}(x+3)=1\] how would I do that?

OpenStudy (anonymous):

start with \[\log(x(x+3))=1\] then \[x(x+3)=10\] and \[x^2+3x-10=0\] etc

OpenStudy (anonymous):

\[e^{-7} e ^{3}e ^{-9}\]

OpenStudy (anonymous):

I think I need to add all of the exponents, but I am not sure what to do after that.

OpenStudy (anonymous):

HELP PLEASE